six less than twice a number is 14. What is the number?
Found 2 solutions by anukamal, Cromlix:
Answer by anukamal(7)

Suppose the number is x
the Equation 2x-6=14
2x=14+6
2x=20
x=10

Answer by Cromlix(4381)

A number = x
2x - 6 = 14
2x = 14 + 6
2x = 20
x = 10
The number is 10.
Hope this helps.
:-)